• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name gaskell

    Log entry time 16:05:13 on November 30, 2008

    Entry number 251503

    keyword=Day Shift Summary

    Day Shift Summary

    8:00 Continue production running on 3He (vertical) from last shift. 11:45 adaql10 stopped again - this kills the online analysis since apparently some scratch disk or smoething is mounted there. 11:55 Beam off for 5 minuites for RICH pedestal run 12:15 Lost ROC23 for HAPPEX daq - don't know how to reboot. 13:00 Bryan called in - ROC is in the next room! Nuts - restarted. 13:50 Analysis worker noticed current was drooping - asked for 1 more uA. 15:25 Another pedestal run for RICH.

    At the end of the shift, Mauro is doing some threshold tests on the RICH.

    Run Events Comment 4058 1 M carryover from last shift 4059 1 M 4060 1 M 4061 1 M 4062 N/A RICH pedestal run 4063 1 M 4064 1 M 4065 1 M 4066 N/A RICH pedestasl run 4067 0.25M Stop to change RICH CRL 4068 100K RICH thresh.changed to 240 4069 100 K RICH threshold to 180